# Lyapunov Dynamics in Entangled Biphoton Spectroscopy

### v1

@@ Abstract
Recent advances in quantum light spectroscopy highlight the potential of using entangled photons as a sensitive probe for many-body dynamics and material correlations. However, a comprehensive theory to explain experimental results remains elusive, primarily due to the complexity of the Hilbert space and the intricate interactions and nonlinearities inherent in material systems. In this work, we introduce a tractable model based on a finite-sized correlation matrix governed by a bilinear bosonic Hamiltonian, enabling efficient simulations through Gaussian-preserving dynamics. We apply this framework to compute the output joint spectral intensity (JSI) and von Neumann entropy of frequency-entangled biphotons, and find close agreement with experimental observations in empty microcavities. Our results reveal the emergence of off-diagonal spectral correlations that can be interpreted as irreversible decay of cavity excitations into the biphoton continua. This approach offers a powerful theoretical tool for interpreting quantum spectroscopic data and paves the way for probing more complex light-matter interactions in materials.
@@ I Introduction
A growing literature in theoretical and experimental quantum light spectroscopy techniques reveals that quantum light can be used as a sensitive tool to probe the electronic and vibrational properties of materials [1, 2, 3, 4, 5, 6, 7, 8, 9]. The ability to perform spectroscopic measurements using low-photon fluxes becomes significant when exploring the energy conversion dynamics in photovoltaic materials, microscopic phenomena in biological systems, and any other such systems where understanding the many-body correlations and fluctuations reveal richer information than classical spectroscopic techniques. Traditionally, quantum spectroscopic experiments look at the changes in entanglement of photons as they interact with the many degrees of freedom of the material considered. Recent progress in this direction employs frequency-entangled biphotons as the probe as seen in Malatesta et al. [10] and Moretti et al. [11] These experiments use a process called Spontaneous Parametric Down Conversion (SPDC) where a pump photon, incident on a nonlinear crystal, generates two channels of photons, conventionally called the signal and idler channels. This process conserves the frequency and momentum of the pump and produces the two channels with a distributed amplitude along a frequency range. We refer to this amplitude as the Joint Spectral Amplitude (JSA). However, in photon-counting experiments, one usually has to deal with a delay between two consecutive coincidence detection events. As a result, the data obtained is effectively a time-integrated probability amplitude over the delay period. This is known as the Joint Spectral Intensity (JSI), which is usually the observable in quantum spectroscopic experiments.
In order to realize a viable platform to conduct these experiments, one needs to fabricate strong light-matter interactions. This is achieved by placing materials within optical microcavities. These not only have the ability to amplify electric fields, but also provide quantized electromagnetic modes (EM) that interact with materials. These quantized modes enable us to probe the dynamics between multiple excitons and provide a comprehensive understanding of the many-body correlations that exist in materials. These EM modes interacting with material excitations form strongly-interacting light-matter hybrid particles called polaritons that offer a testbed to realize a plethora of exotic physics. These include Bose-Einstein Condensation of polaritons [12, 13, 14], superfluidity and vortex formations [15, 16], nonlinear effects such as optical bistability [17, 18], four-wave mixing [19], and soliton formation [20, 21]. Moreover, in specially engineered photonic lattices, polaritons can exhibit topological protection, similar to the quantum Hall effect or topological insulator phases [22, 23]. In two-dimensional planar materials such as and graphite, phonon polaritons can exhibit hyperbolic dispersion [24, 25, 26], enabling deep subwavelength confinement. Since these polaritons represent strong light-matter interactions, they form the perfect spectroscopic tool to probe many-body correlations using single- and biphoton sources.
Interpreting experimental results requires a robust theoretical framework in quantum spectroscopy. Although ample theoretical development is available in the literature, they fall short in reproducing experimental observations and accurately explaining the physics behind polariton-induced scattering. Bittner et al. [27] showed how radiative cascade effects and biphoton scattering among molecular dimers lead to a generation in the output entanglement entropy of biphoton states. In this, they treat the dimer as coupled excitons with anharmonic interactions embedded within a continuum of photons and show, using diagrammatic expansions, that the coupling strength and repulsions can be correlated to the output entropy. In addition, Li et al. [28, 29] also laid a similar framework to establish photon-photon correlations using the Dicke model and the input-output theory [30, 31] in a Hong-Ou-Mandel apparatus [32]. In this apparatus, only one of the photons of the biphotons produced from the SPDC process interacts with a resonant medium. As this photon passes through the medium, a transmission function is introduced that is multiplied by the input to give the output photon state. All many-body dynamics of the resonant medium is then traced back by studying the pole structure of this function and analyzing the dynamics of the resulting polariton branches. While these developments provide detailed frameworks for describing many interesting and non-linear photophysics, their numerical simulations become quite challenging owing to an exponentially growing Hilbert space from the many biphoton modes. For example, for a biphoton state interacting with a cavity-polariton under the Jaynes-Cummings Hamiltonian, the Hilbert space is of the order of where is the number of signal/idler photon modes. Moreover, it becomes imperative to describe the mapping between the input and output modes of the system when attempting to reconcile with tradional experimental practices. In particular, the development of a numerically-tractable framework that directly connects the input and output JSIs can prove to be a fruitful foundation upon which more complicated photophysics can be built.
To address these gaps, we present a simple model based on the dynamics of correlation matrices. We consider the governing Hamiltonian to comprise bilinear bosonic operators. Consequently, the time-evolution unitary operator, , becomes a gaussian preserving map. This allows us to translate our dynamics into the first moments and covariance matrices of the problem. As a result, we can perform numerical simulations using finite matrices and arrive at convergent solutions to the output JSI. This can be done by solving the resultant Lyapunov equations of the covariance matrix and extract the output correlations between the signal and idler photons. These correlations represent the output JSA which can then be compared to the input.
Additionally, we use Møller operators to connect the input, interacting, and output variables of our experiment, eliminating the need for discrete time gating and yielding a framework that is commensurate with experimental practice. These operators provide an elegant method to connect the input and output covariance matrices that are mediated by the Lyapunov equations. Other observables, like the purity of the output biphoton state, can also be seamlessly extracted from the Wigner function expressed in terms of the covariance matrix. This framework also allows us to easily extend our study to the Tavis-Cummings limit, enabling a deeper probe of material systems. We demonstrate that with these considerations, our model predicts an output JSI that bears close resemblance with experimental observations of biphotons interacting with an empty microcavity.
@@ II.1 Hamiltonian
where are the cavity modes, are the biphoton signal/idler operators, and the are operators of the material modes that are assumed to be bosonic. The first 3 terms of Eq. (II.1) are the free Hamiltonians of the cavity, signal/idler photons and the material respectively. The fourth term describes the bilinear coupling between the signal/idler modes with the photons and the last term describes the coupling between the cavity and material modes. The coupling terms obey the first Markov and rotating wave approximations. The former assumes a uniform coupling strength and the latter removes rapid counter-rotating terms.
@@
where is the joint spectral amplitude of the input photons. This two-dimensional correlation function quantifies the degree of entanglement between the signal and idler photon states.
Fig. 1 shows an example of an experimental input and output biphoton JSI for empty microcavities. The input JSA represents a spectrally entangled state that is generated in a Type-I -Barium Borate (BBO) crystal phase-matched for SPDC close to the degeneracy at the pump wavelength of nm [10]. This dictates the signal/idler dispersion in Eq. (II.1). Several different input states have also been experimentally produced as demonstrated in Moretti et al. [11] Similarly, the output state can be defined as,
The subscripts ”in” and ”out” in the operators act on the input and output vacuum respectively. According to the input-output formalism [30, 31], the output and the input operators are related to the coupled cavity mode as,
@@ II.2 Governing Equation of the Model
The Hamiltonian in Eq. (II.1) can be numerically intractable with current computational capabilities. In order to make this problem solvable, we note that all the operators in the Hamiltonian are bilinear. As a result, the unitary operator is a gaussian-preserving map, allowing us to reflect the dynamics of our system in the first moments and covariances [34]. We can also use Wick’s theorem to contract higher-order moments into these first- and second-order moments. This is useful since we can perform our dynamics with finite-matrix theory without dealing with an exponentially growing Hilbert space. The equation of motion for the operators can then be written in a linearized form as,
@@
where Eq. (13) has been used. Eq. (17) is known as the Sylvester differential equation and its stationary form is the well-known Lyapunov equation used in the study of chaos [35]. Solving the Sylvester equation allows us to understand the evolution of signal-idler correlations as they interact with a microcavity. In solving this, we must ensure that it concurs with the boundary conditions of the correlation matrix. To establish this, we convert it into its respective input and output forms as follows:
@@
where is the correlation matrix of the output. We take so that is the integration of correlations over all time. Since the biphotons are taken to be non-interacting at asymptotic times and interacting at finite times, this time-integrated quantity is effective in capturing the dynamics of the experiment. It evolves under the full Hamiltonian and is useful in setting up our boundary conditions. Fixing , we obtain,
@@
Let us assign . This matrix takes the form of Møller operators as defined in Eqs. (8) - (10). Eq. (27) physically says that the input modes are propagated forward in time and the output modes are propagated backward in time under the interacting Hamiltonian. In doing so, both modes must meet at a common point in time, , to maintain continuity. This forms the essence in which we connect the input and output Heisenberg operators. Since the covariance matrix is defined as,
@@
This model is also robust to the number of material degrees of freedom present in the cavity as long as the Hamiltonian in Eq. (II.1) generates a gaussian-preserving map. This means that it can be applied equally well to monomeric, dimeric, and other polymeric systems under the two-level approximation. An illustration of such an experimental setup and the theoretical boundary conditions that we described earlier are shown in Figs. 2(a) and 2(b).
@@ II.4 Entanglement Entropy
To derive the purity of the state, we Weyl transform the problem and express the corresponding Wigner function of our multi-mode state in terms of the covariance matrix as descrobed in Ref [34]. This takes the form,
@@ III.2 Experimental Initial JSI
To benchmark the model, we take an experimental input JSI and observe the dynamics. In Fig. 4, we see that the JSI shifts towards higher idler and lower signal wavelengths. Since the system is set up such that the cavity and material dispersions are nearly in resonance with the idler wavelength (), these energies are preferentially absorbed by the cavity and material with a higher probability as compared to the signal photons. As a result, we see a shift in the JSI towards higher idler wavelengths (lower frequencies). For higher coupling strengths, this process occurs at a much faster time scale because of an increase in the Rabi frequency. As a result, for asymptotic output times, the JSI tends to saturate toward higher(lower) idler(signal) wavelengths for higher values of the coupling . The reciprocal behavior between the signal and idler shifts is due to the conservation of energy. We also see a squeezing of the map indicating the filtering effect of cavities. This result is in close agreement with the experimental study conducted in empty microcavities [10] as seen in Fig. 1
@@
Bittner et al. [27] also proved that by measuring the change in the photon entanglement entropy, one can find a direct measure of material correlations. To quantify this, we study its von Neumann entropy given in Eqs. (33) and (34). In Fig. 6, we see that the output entanglement entropy decreases monotonically as a function of the coupling strength, . This occurs because the relatively preferential absorption of the idler photons by the cavity and material modes degrades the correlations between the signal and idler photons. For dimers, the rate of decrease is higher than that for monomers, as seen by the red curve. This is because excitations are able to explore a higher Hilbert space and delocalize into these modes faster in the same amount of time. An intriguing aspect of this plot is that the entropy increases after a certain value of the coupling strength. This points toward a revival of strong correlations in the output signal and idler photons.
@@
Fig. 4 also reveals the emergence of peaks along the off-diagonal, indicating an amplification of signal-idler correlations between specific frequency modes. This behavior can be understood by recognizing that our system comprises two discrete modes—the cavity and material excitations—and two continua of states: the signal and idler modes. To capture the dynamics of a discrete mode coupled to a continuum, one typically discretizes the continua, computes the transition amplitudes, and then takes the continuum limit by letting the discretization spacing approach zero. This procedure enables the calculation of physically relevant quantities, such as level shifts, decay rates, cross sections, and more. A detailed account of this formalism is provided in Cohen-Tannoudji et al. [36] When a discrete mode couples to multiple discrete states that are well isolated from other states in the system, the resulting transition amplitude becomes a superposition of Rabi oscillations with varying frequencies and amplitudes. In the continuum limit, these superpositions lead to an effective irreversible decay of the discrete mode into the continuum. The resulting overlap of the discrete state with the quasi-continuum forms a Lorentzian lineshape centered at the discrete state energy with a width , where denotes the decay rate of the mode.
@@
In our setup, when the discrete cavity mode couples to the signal-idler continua, the cavity excitations decay equally into both continua for the same values of signal-idler frequencies. Because the initial correlations are antisymmetric with respect to the signal and idler, and lie along the diagonal, this decay process leads to an emergence of off-diagonal correlations, manifesting as peaks in Fig. 4. The JSI plotted here is obtained by tracing out the signal-idler submatrix from the final correlation matrix of the complete system and thus reflects the excitations that have decayed into the quasi-continuum.
@@
Additionally, if the material were coupled to the continua in such a way that the ratio of its couplings did not lie in limiting regimes, the resulting output states would exhibit features characteristic of Fano-type asymmetric lineshapes. The trends observed above highlight the capability of photon entanglement to serve as a sensitive probe for exploring many-body interactions and correlations in quantum spectroscopy.
The simplified structure of our model, while not capturing all nonlinear and higher-order effects, enables analytical tractability and provides a lucid method to connect the input and output variables. It successfully reproduces key experimental features and serves as a useful baseline for interpreting the role of entanglement in spectroscopic signals. However, this same formalism necessarily omits certain processes, such as pure dephasing, typically modeled via -type jump operators in the Lindbladian. Stated succinctly, the assumptions of bilinearity and linear coupling prevent the model from capturing non-Markovian dynamics, crosstalk, and higher-order excitations. These effects often arise in experimental systems due to energetic disorder, field inhomogeneity, and cascaded transitions [29]. However, our formalism provides a robust starting point for systematically extending the model to include such interactions. This forms the basis of our future investigation.

### v2

# Quantum Spectroscopy with Biphotons: Lyapunov-Based Input–Output Dynamics

@@ Abstract
We develop a Lyapunov-based framework to model the evolution of entangled biphotons interacting with cavity and material modes. Using Gaussian-preserving dynamics and Møller operators, we map input joint spectral amplitudes to experimentally measurable joint spectral intensities. Our model reproduces key features of observed spectra and reveals off-diagonal correlations arising from cavity decay, providing a scalable and tractable tool for quantum spectroscopic analysis.
@@ I Introduction
Recent advances in quantum light spectroscopy have revealed that quantum correlations in light can serve as powerful probes of electronic and vibrational dynamics in condensed-phase systems [1, 2, 3, 4, 5, 6, 7, 8, 9]. By operating in the few-photon regime, these techniques minimize sample perturbation and access information encoded in many-body correlations—features that are often masked in classical measurements. This sensitivity becomes essential when probing ultrafast energy conversion processes in photovoltaic materials or fragile quantum coherence in biological and molecular systems.
A central thrust in this field now focuses on how entangled photons evolve upon interacting with complex matter. This interaction imprints signatures of the material’s internal correlations onto the joint properties of the light field [10, 11, 12, 13, 14]. Specifically, spectroscopic protocols that employ time-frequency entangled photon pairs generated via Spontaneous Parametric Down Conversion (SPDC) offer new pathways to interrogate material dynamics with high spectral and temporal resolution [15, 16]. In SPDC, a single pump photon generates a correlated pair—signal and idler—distributed over a continuum of frequencies while conserving energy and momentum. The two-photon state is characterized by the joint spectral amplitude (JSA), which encodes both amplitude and phase correlations between the pair.
In practice, time-resolved coincidence counting introduces integration over the detection delay, effectively collapsing the measurement to the joint spectral intensity (JSI)—the squared modulus of the JSA. This object, accessible in experiment, captures the frequency-domain correlations shaped by both the quantum state of the probe and its interaction with the material system. As we demonstrate below, the structure of the JSI encodes key features of the underlying light-matter dynamics and provides a versatile platform for extracting coherence and correlation information beyond the reach of classical spectroscopy.
To overcome the intrinsically low probability of light-matter interactions at the single-photon level, quantum spectroscopic protocols often require operation in the strong coupling regime. This regime is realized by embedding materials within optical microcavities, which enhance the interaction strength by both amplifying the local electric fields and quantizing the electromagnetic (EM) modes that couple to material excitations. These discrete EM modes not only mediate efficient coupling between light and matter, but also enable access to dynamics involving multiple excitons, thereby revealing the structure of many-body correlations that govern material response.
When the coupling strength exceeds the dissipative losses, light and matter hybridize to form polaritons—quasiparticles that inherit characteristics of both photons and material excitations. Polaritons serve as a robust platform for exploring collective quantum phenomena, including Bose-Einstein condensation [17, 18, 19], superfluid transport and vortex formation [20, 21], and nonlinear optical behavior such as bistability [22, 23], four-wave mixing [24], and soliton propagation [25, 26]. In suitably engineered photonic lattices, polaritons can acquire topological properties, exhibiting chiral edge modes and disorder-immune transport analogous to the quantum Hall effect and topological insulators [27, 28]. Furthermore, in two-dimensional van der Waals materials such as and graphite, phonon-polaritons exhibit hyperbolic dispersion [29, 30, 31], enabling deep subwavelength confinement and directional energy flow. These strongly coupled light-matter systems present an ideal testbed for quantum spectroscopy. By probing their response with single- and biphoton sources, one can extract signatures of many-body interactions, quantum correlations, and nonclassical dynamics that remain hidden in conventional optical measurements.
Interpreting experimental results in quantum spectroscopy demands a rigorous theoretical framework, and several such approaches have recently emerged. Bittner [10] demonstrated that biphoton scattering and radiative cascade processes in coupled molecular dimers can generate entanglement entropy in the outgoing biphoton state. Modeling the dimer as a pair of anharmonically interacting excitons embedded in a photonic continuum, they employed diagrammatic techniques to show that the output entropy correlates with both the interaction strength and the repulsive coupling. Complementary work by Li and coworkers [11, 12] developed a related framework using the Dicke model and input-output theory [32, 33] to analyze photon-photon correlations in a Hong-Ou-Mandel (HOM) configuration [34]. In this scheme, only one photon from a time-frequency entangled pair interacts with a resonant medium, and its output state is shaped by a transmission function that encodes the system’s spectral response. The poles of this function define the underlying polariton branches and reveal the many-body dynamics of the medium.
While these models successfully capture nonlinear photophysics and many-body correlations, they often involve Hilbert spaces that scale exponentially with photon number and mode discretization. For instance, modeling a biphoton interacting with a cavity polariton via the Jaynes-Cummings Hamiltonian yields a Hilbert space dimension of order , where is the number of signal and idler photon modes. This exponential growth poses significant challenges for numerical simulation and data interpretation. Moreover, directly mapping the input JSA to the output JSI—the experimentally accessible observable—remains a nontrivial task in many of these models. Developing a reduced, yet physically accurate, framework that captures this mapping while remaining computationally tractable would provide a robust foundation for extending quantum spectroscopy to more complex and strongly correlated systems.
We employ Møller operators to connect the input, interacting, and output states of the system, thereby eliminating the need for discrete-time gating and establishing a framework directly aligned with experimental practice. These operators provide a natural mechanism for linking the input and output covariance matrices, with the transformation governed by Lyapunov equations [35]. This formalism enables the direct extraction of key observables, such as the purity of the output biphoton state, from the Wigner function expressed in terms of the covariance matrix. Importantly, the same framework generalizes naturally to the Tavis-Cummings limit, allowing for systematic exploration of collective light-matter interactions in many-body systems. Within this approach, we show that the predicted output JSI reproduces key features observed in experiments involving biphoton interactions with an empty microcavity, underscoring the consistency and applicability of our model.
@@ II.1 Hamiltonian
where are the cavity modes, are the biphoton signal/idler operators, and the are operators of the material modes that are assumed to be bosonic. The first 3 terms of Eq. (II.1) are the free Hamiltonians of the cavity, signal/idler photons, and the material, respectively. The fourth term describes the bilinear coupling between the signal/idler modes of the photons, and the last term describes the coupling between the cavity and material modes. The coupling terms obey the first Markov and rotating wave approximations. The former assumes a uniform coupling strength, and the latter removes rapid counter-rotating terms.
@@
where is the joint spectral amplitude (JSA) of the input photons. This two-dimensional correlation function quantifies the degree of entanglement between the signal and idler photon states.
Fig. 1 shows an example of an experimental input and output biphoton JSI for empty microcavities. The input JSA represents a spectrally entangled state that is generated in a Type-I -Barium Borate (BBO) crystal phase-matched for SPDC with the pump wavelength of nm [15]. This dictates the signal/idler dispersion in Eq. (II.1). Several different input states have also been experimentally produced as demonstrated in Moretti et al. [16]. Similarly, the output state can be defined as,
The subscripts “in” and “out” in the operators act on the input and output vacuum respectively. According to the input-output formalism [32, 33], the output and the input operators are related to the coupled cavity mode as,
@@ II.2 Governing Equation of the Model
The Hamiltonian in Eq. (II.1) can be numerically intractable with current computational capabilities. To make this problem solvable, we note that all the operators in the Hamiltonian are bilinear. As a result, the unitary operator is a Gaussian-preserving map, allowing us to reflect the dynamics of our system in the first moments and covariances [37]. We can also use Wick’s theorem to contract higher-order moments into these first- and second-order moments. This is useful since we can perform our dynamics with finite-matrix theory without dealing with an exponentially growing Hilbert space. The equation of motion for the operators can then be written in a linearized form as,
@@
where Eq. (13) has been used. Eq. (17) is known as the Sylvester differential equation, and its stationary form is the well-known Lyapunov equation used in the study of chaos [35]. Solving the Sylvester equation allows us to understand the evolution of signal-idler correlations as they interact with a microcavity. In solving this, we must ensure that it is consistent with the boundary conditions of the correlation matrix. To establish this, we convert it into its respective input and output forms as follows:
@@
where is the correlation matrix of the output. We take so that is the integration of correlations over all time. Since the biphotons are taken to be non-interacting at asymptotic times and interacting at finite times, this time-integrated quantity effectively captures the experiment’s dynamics. It evolves under the full Hamiltonian and is useful in setting up our boundary conditions. Fixing , we obtain,
@@
Let us assign . This matrix takes the form of Møller operators as defined in Eqs. (8) - (10). Eq. (27) physically says that the input modes are propagated forward in time and the output modes are propagated backward in time under the interacting Hamiltonian. In doing so, both modes must meet at a common point in time, , to maintain continuity. This forms the essence of the connection between the input and output Heisenberg operators. Since the covariance matrix is defined as,
@@
This model is also robust to the number of material degrees of freedom present in the cavity as long as the Hamiltonian in Eq. (II.1) generates a Gaussian-preserving map. This means that it can be applied equally well to monomeric, dimeric, and other polymeric systems under the two-level approximation. An illustration of such an experimental setup and the theoretical boundary conditions that we described earlier is shown in Figs. 2(a) and 2(b).
@@ II.4 Entanglement Entropy
To derive the purity of the state, we Weyl-transform the problem and express the corresponding Wigner function of our multi-mode state in terms of the covariance matrix as described in Ref [37]. This takes the form,
@@ III.2 Experimental Initial JSI
To benchmark the model, we take an experimental input JSI and estimate the dynamics. In Fig. 4, we see that the JSI shifts towards higher idler and lower signal wavelengths. Since the system is set up such that the cavity and material dispersions are nearly in resonance with the idler wavelength (), these energies are preferentially absorbed by the cavity and material with a higher probability as compared to the signal photons. As a result, we see a shift in the JSI towards higher idler wavelengths (lower frequencies). For higher coupling strengths, this process occurs at a much faster time scale because of an increase in the Rabi frequency. As a result, for asymptotic output times, the JSI tends to saturate toward higher(lower) idler(signal) wavelengths for higher values of the coupling . The reciprocal behavior between the signal and idler shifts is due to the conservation of energy. We also see a squeezing of the map, indicating the filtering effect of cavities. This result is in close agreement with the experimental study conducted in empty microcavities [15] as seen in Fig. 1
@@
Bittner et al. [10] also proved that by measuring the change in the photon entanglement entropy, one can find a direct measure of material correlations. To quantify this, we study its von Neumann entropy given in Eqs. (33) and (34). In Fig. 6, we see that the output entanglement entropy decreases monotonically as a function of the coupling strength, . This occurs because the relatively preferential absorption of the idler photons by the cavity and material modes degrades the correlations between the signal and idler photons. For dimers, the rate of decrease is higher than that for monomers, as seen by the red curve. This is because excitations can explore a higher Hilbert space and delocalize into these modes faster in the same amount of time. An intriguing aspect of this plot is that the entropy increases after a certain value of the coupling strength. This points toward a revival of strong correlations in the output signal and idler photons.
@@
Fig. 4 also reveals the emergence of peaks along the off-diagonal, indicating an amplification of signal-idler correlations between specific frequency modes. This behavior can be understood by recognizing that our system comprises two discrete modes—the cavity and material excitations—and two continua of states: the signal and idler modes. To capture the dynamics of a discrete mode coupled to a continuum, one typically discretizes the continuum, computes the transition amplitudes, and then takes the continuum limit by letting the discretization spacing approach zero. This procedure enables the calculation of physically relevant quantities, such as level shifts, decay rates, cross sections, and more. A detailed account of this formalism is provided in Cohen-Tannoudji et al. [38]. When a discrete mode couples to multiple discrete states that are well isolated from other states in the system, the resulting transition amplitude becomes a superposition of Rabi oscillations with varying frequencies and amplitudes. In the continuum limit, these superpositions lead to an effective irreversible decay of the discrete mode into the continuum, forming a quasi-continuum. The resulting overlap of the discrete state with the quasi-continuum forms a Lorentzian lineshape centered at the discrete state energy with a width , where denotes the decay rate of the mode.
@@
In our setup, when the discrete cavity mode couples to the signal-idler continua, the cavity excitations decay equally into both continua for the same values of signal-idler frequencies. Since the initial correlations are antisymmetric with respect to the signal and idler, and lie along the diagonal, this decay process leads to the emergence of off-diagonal correlations, manifesting as peaks in Fig. 4. The JSI plotted here is obtained by tracing out the signal-idler submatrix from the final correlation matrix of the complete system, and thus reflects the excitations that have decayed into the quasi-continuum.
@@
Additionally, if the material were coupled to the continua in such a way that the ratio of its couplings did not lie in limiting regimes, the resulting output states would exhibit features characteristic of Fano-type lineshapes. In such a scenario, the JSA lineshapes developed in the output would be asymmetric about the cavity resonance. When applied to both the signal and idler channels, and depending on the detuning between the cavity and material, we expect a superposition of each of these asymmetric lineshapes to appear in the off-diagonal region of the output JSI. These observations would then be characteristic of many-body interactions occuring in our material-cavity system. The trends observed above highlight the capability of photon entanglement to serve as a sensitive probe for exploring many-body interactions and correlations in quantum spectroscopy.
The simplified structure of our model offers significant analytical tractability and a framework to connect input and output observables. Despite its simplicity, it successfully reproduces the experimentally obtained JSI, and establishes a practical baseline for interpreting the role of entanglement in quantum spectroscopic signals. That said, the formalism inherently excludes certain physical processes, such as pure dephasing, which are typically modeled by -type jump operators within the Lindbladian. By assuming the Hamiltonian to be in a bilinear form, we preclude the description of non-Markovian dynamics, inter-mode crosstalk, and multi-photon interactions. These effects may arise from energetic disorder, spatial inhomogeneity, or cascaded transitions [12]. We plan to use the present formalism as a robust foundation and incorporate these complexities in our future investigation.